OpenStudy (anonymous):

Find the real solutions of the equation: (6x+5)^(2)-6(6x+5)+9=0 Thx!! ;)

OpenStudy (experimentx):

first of all, simplify it ax^2+bx+c=0, and you have quadratic equation whose roots is given by http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (paxpolaris):

Let 6x +5 = A \[\large A^2-6A+9=0\]\[\large \implies \left( A-3 \right)^2=0\] \[A=3\]

OpenStudy (paxpolaris):

\[\therefore 6x+5 =3\]\[\Large x=-\frac13\]